OpenStudy (anonymous):

Prove that cos(2s) = cos^2s - sin^2s

OpenStudy (anonymous):

\[\cos (2 s) = \cos (s + s)\]

OpenStudy (anonymous):

Nw use ..... cos (x+y) = cosx cosy - sinx siny.hope this helps

OpenStudy (anonymous):

\[\cos(s+s) = \cos (s) . \cos(s) - \sin(s).\sin(s)\]

OpenStudy (anonymous):

= \[\cos^2 (s) - \sin^2 (s)\]

OpenStudy (anonymous):

ah thank you

OpenStudy (anonymous):

Welcome...)